OpenStudy (anonymous):

What point is the intersection of the graphs of x2 + y2 = 25 and y = x + 7?

OpenStudy (campbell_st):

use substitution x^2 + y^2 =25 is a circle with a centre at the origin and radius 5 y = x + 7 is a straight line with a y intercept at y = 7 then x^2 + (x + 7)^2 = 25 now expand and simplify x^2 + x^2 + 14x + 49 = 25 2x^2 + 14x + 24 = 0 2(x^2 + 7x + 12) = 0 which can be factorised to 2(x +3)(x+4) = 0 I'll leave you to find the 2 points where the curves intersect.